2x+4y=-10
-3x-2y=-1

Answer by jim_thompson5910(35256)

First, double the second equation

2x+4y=-10
-3x-2y=-1


2x+4y=-10
-6x-4y=-2


Then add the two equations to get -4x=-12

Solve for x to get x = 3


Then plug that into either equation and solve for y

2x+4y=-10

2(3)+4y=-10

6+4y=-10

4y=-10-6

4y=-16

y=-4


Answers are x = 3 and y = -4